Owl shift 99/29/05

Nickolas Solomey (solomey@hep.uchicago.edu)
Sat, 29 May 1999 06:56:39 -0500

Shift Summary for owl shift, 99/29/05

On Shift: Ashkan and Nick, plus almost everyone else at the beginning
then Rick, Sasha, Hogan, Migeal and Theo decided to stay and make sure
we stayed awake. Greg Bock kept phone from the MCR to make sure we
were still here and tried offering us more intensity for no extra
charge.

Hours of beam delivered at all intensities: 4.5
Hours of beam delivered at nominal intensity: 0
Hours of data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 4.5

Accounting for lost beam time: 2.5 hours. No beam and 5 min. Access

========================================================================
Runs ended this shift:
Run 13198-13208 Almost all runs for DC high voltage studies with muons.

Tapes brought to FCC this shift: none

========================================================================
Description of Shift:

Came on shift, hall is interlocked and some low intensity beam
muons were being delivered to the KTeV experimental area. Muon data
taking and DC studies in progress.

MCRM> 0:28 Tev beam being held off due to an HFU at F2 being replaced.
MCRM> 0:56 Beam back, we continue just doing muons to study dc and csi.

2:15-2:20 Hogan and Nick Made access to reset the repeater by the laser
crate, which hanged up stream 5 of the readout.

3:30 Did gas check list, this was not done on the previous nights owl
shift, so if you are on shift don't forget to do the gas check list.

4:00-5:13 Tried to use some of the monitoring code that shift
people are to display, found that ytf, fpedcal, hypmon, and sciamon don't
work, but everything else seemed to have histograms with data in
them which does not necessaryly mean they are working just that their
histograms are filled and present experts will have to check their
own monitoring processes, i.e. most had frowns.

MCRM> 5:13 Beam delayed due a power supply failure at MI-10.
MCRM> 5:19 Contacting experts for power supply failure.
MCRM> 5:39 Power supply expert on route (that means they are not on
vacation this weekend!: commentary by ns)

5:57 Sasha and Migeal make access to raise DC threshold 50 mV to study
DC performance parameters when muons return.

6:25 Greg Bock called from control room, said he was going home
but there was a good chance we could see more beam, "the expert
may arrive and solve the problem trivally or on the other hand there is a
chance it is a more serious problem".

6:30 Ashkan and Theo make access for slow daq work.

6:50 Next shift arrived early, so thats all for me until next weekend.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey